# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
This section contains the proof of Theorem B of the abstract (see Corollary 4.2). Recall from Definition 2.4 and Remark 2.5 that the notion of a faithful extension is a generalization of the notion of a conservative extension. The notion of a $\Sigma_{\omega}$-satisfaction class was defined in Definition 2.7(b).
4.1. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathrm{ZFC}$ such that $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},~{}\mathrm{faith}}\mathcal{N}$, $\mathcal{N}$ fixes $\omega^{\mathcal{M}},$ and $\mathcal{N}$ is taller than $\mathcal{M}$. Then:
(a) There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$. [^16] As pointed out in Remark 4.5, both case of the dichotomy can be realized.
(b) In particular, there is a $\Sigma_{\omega}$-satisfaction class $S$ $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof [^17] The proof strategy of Theorem 4.1 is a variant of the proofs of the following results: [En-2, Theorem 1.5], [En-3, Theorem 3.3], [EH, Theorem 2.1], [En-6, Theorem 2.1.3], and [En-7, Lemma 2.19]. A variant of the same strategy is used in the proof of Theorem 4.5. The origins of the strategy can be traced to Kaufmann’s refinement [Kay, Lemma 3.2] of a Skolem hull argument due to Keisler and Silver [KS, Theorem 2.1].. Fix some $\lambda\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ and some $\mathcal{N}$-ordinal $\beta>\lambda$. Note that:
@@
$K_{m}:=\{a\in\mathrm{V}_{\beta}:\mathcal{N}\models a\in\mathrm{Def}($$V$,${}_{\beta},\in,\vartriangleleft,\lambda,m)\},$
where $x\in\mathrm{Def}($$V$,${}_{\beta},\in,\vartriangleleft,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in the structure (((($V$,${}_{\beta},\in,\vartriangleleft,\lambda,m).$
Note that the only allowable parameters used in a definition of $x$ are $\lambda$ and $m$. Thus, intuitively speaking, $K_{m}$ consists of elements $a$ of $N_{\beta}$ such that within $\mathcal{N}$, $s$ is first order definable in $\left(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m\right)$. Notice that since we are not assuming that $\mathcal{M}$ is $\omega$-standard, an element $a\in K_{m}$ need not be definable in the real world in the structure $\left(\mathcal{N}_{\beta},\vartriangleleft,\lambda,m\right).$ Next we move outside of $\mathcal{N}$ and define $K$ as follows:
@@
(2) $\ \ \mathcal{M}\subsetneq\mathcal{K}\preceq\mathcal{N}_{\beta}$.
@@
(3) $\ \ \mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{K}\preceq\mathcal{N}_{\beta}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$.
@@
Clearly $O^{\ast}$ is nonempty since $\lambda\in O$ We now consider the following two cases. As we shall see, Case I leads to the conclusion of the theorem, and Case II is impossible.
@@
(4) $\ \ \mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cof}}\mathcal{K}_{\gamma_{0}}.$
@@
$\Phi_{m}:=\{x\in M:\mathcal{N}\models$ $x$ is (the code of) a formula $\varphi(c,c_{m})$ such that $\varphi(c_{\lambda},c_{m})\in s_{\beta}\}\mathrm{.}$
So intuitively speaking, $\Phi$ is the result of replacing $c_{\lambda}$ by $c$ (where $c$ is a fresh constant) in the sentences in the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) whose constants are in $\{c_{\lambda}\}\cup\{c_{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ might be nonstandard formulae.
@@
$\Gamma:=\left\{t(c,c_{m})\in M:t(c,c_{m})\in\Phi\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,c_{m})>c_{\theta}\right)\in\Phi\right\},$
where $t$ is a definable term in the language $\mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{c_{m}:m\in M\}.$ Officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,c_{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
$(i)\qquad\left[\exists!x\varphi(c,c_{m},x)\right]\in\Phi.$
$(ii)\qquad\left[\forall x\left(\varphi(c,c_{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
$(iii)\qquad\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow c_{\theta}\in x\right)\right]\in\Phi.$
@@
$\psi:=\forall t\left(t\in\Gamma\longrightarrow\left[\exists t^{\prime}\in\Gamma\wedge\left((t^{\prime}\in t)\in\Phi\right)\right]\right).$
The veracity of the dependent choice scheme in $\mathcal{M}$ (see part (e) of Definition 2.1), together with the facts that $\Phi$ is $\mathcal{M}$-amenable and $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$ make it clear that there is a sequence $s=\left\langle t_{n}:n\in\omega^{\mathcal{M}}\right\rangle$ in $\mathcal{M}$ such that:
@@
$\mathcal{N}\models\left\langle t_{n}^{(\mathrm{V}_{\beta},\in)}(c_{\lambda}):n\in\omega\right\rangle$ is an infinite decreasing sequence of ordinals.
In the above $t_{n}(c_{\lambda})$ is the term obtained by replacing $c$ with $c_{\lambda}$ in $t_{n},$ and $t_{n}^{(\mathrm{V}_{\beta},\in)}(c_{\lambda})$ is the interpretation of $t_{n}(c_{\lambda})$ in $(\mathrm{V}_{\beta},\in)$. $\square$
4.2. Corollary. If $\mathcal{M}\models\mathrm{ZFC}$, $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ fixes $\omega^{\mathcal{M}},$ then $\mathcal{M}$ is cofinal in $\mathcal{N}$. Thus a conservative elementary extension of $\mathcal{M}$ that fixes $\omega^{\mathcal{M}}$ is a cofinal extension.
Proof. If $\mathcal{M}\models\mathrm{ZFC}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ is not cofinal in $\mathcal{N}$, then $\mathcal{N}$ is taller than $\mathcal{M}$. But then by part (b) of Theorem 4.1, there is an $\mathcal{M}$-definable $\Sigma_{\omega}$-satisfaction class over $\mathcal{M}$, which contradicts Tarski’s Undefinability of Truth Theorem. $\square$
In Theorem 4.4 below we show that the conclusion of Theorem 4.1 holds if $\mathrm{ZFC}$ is weakened to $\mathrm{ZF,}$ but the assumption that $\mathcal{N}$ fixes $\omega^{\mathcal{M}}$ is strengthened to $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}$. We first establish a lemma that will come handy in the proof of Theorem 4.4.
4.3. Lemma Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then $\mathcal{M}\subsetneq_{\mathrm{rank}}\mathcal{N}$.
@@
(2) $\ \ \mathrm{Ext}_{\mathcal{M}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{M}}(a),$ and
@@
So by (1) and (2),
@@
(5) $\ \ \mathrm{Ext}_{\mathcal{N}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)$,
which by Extensionality makes it clear that $b^{\prime}=b.$square
  • The proof of Theorem 4.4 below is a variant of the proof of Theorem 4.1, and will strike the reader who has worked through the proof of Theorem 4.1 as repetitious, but the proof is a bit more involved here since we do not have access to the axiom of choice in $\mathcal{M}$ or in $\mathcal{N}$. Note that in Theorem 4.4 $\mathcal{N}$ is assumed to end extend $\mathcal{M}$, whereas in Theorem 4.1 $\mathcal{N}$ is only assumed to fix $\omega^{\mathcal{M}}.$ Of course Theorem 4.4 follows from Theorem 4.1 if $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZFC}$.
4.4. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then:
(a) There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$.
(b) In particular, there is a $\Sigma_{\omega}$-satisfaction class $S$ $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof. Assume $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the theorem. By Lemma 4.3 we can assume that $\mathcal{N}$ rank extends $\mathcal{M}$. Fix some $\lambda\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ and some ordinal $\beta>\lambda$. For each $m\in M$, we can define the following set within $\mathcal{N}$:
@@
where $x\in\mathrm{Def}($$V$,${}_{\beta},\in,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in (((($V$,${}_{\beta},\in)$ using at most the parameters $\lambda$ and $m.$
@@
Let $O^{\ast}=O\backslash\mathrm{Ord}^{\mathcal{M}}$. Clearly $O^{\ast}$ is nonempty since $\lambda\in O^{\ast}.$ We now consider two cases:
(I) $\ \ O^{\ast}$ has a least element (under $\in^{\mathcal{N}}).$
(II) $\ \ O^{\ast}$ has no least element.
Case I. Let $\gamma_{0}$ be the least element of $O^{\ast}.$ It is easy to see that $\gamma_{0}$ is a limit ordinal in $\mathcal{N}$. We claim that
@@
To verify the claim, by the Tarski criterion of elementarity, it suffices to show that if $\mathcal{N}(\gamma_{0})\models\exists y\varphi(y,m)$ for some $m\in M$, then there is some $m_{0}\in M$ such that $\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$ Let $\gamma_{m}$ be the first ordinal below $\gamma_{0}$ such that:
@@
So $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameters $\gamma_{0}$ and $m$, and since $\gamma_{0}\in O$ by assumption, this shows that $\gamma_{m}$ is definable in $\mathcal{N}(\beta)$ with parameter $m.$ So $\gamma_{m}\in O.$ By the fact that $\gamma_{0}$ was chosen to be the least element of $O^{\ast}$, this shows that $\gamma_{m}\in\mathrm{Ord}^{\mathcal{M}},$ which makes it clear that there is some $m_{0}\in M$ such that:
@@
This concludes the verification of $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. To finish the proof of Theorem 4.4 it suffices to show that Case II is impossible. Suppose to the contrary that $\mathrm{Ord}^{\mathcal{K}}\backslash\mathrm{Ord}^{\mathcal{M}}$ has no least element. Within $\mathcal{N}$ let $s_{\beta}$ be the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$. For each $m\in M$ let
$\Phi_{m}:=\{\varphi(c,c_{m})\in M:\mathcal{N}\models$ $\varphi(c_{\lambda},c_{m})\in s_{\beta}\}\mathrm{,}$
and let
@@
Note that $\Phi\subseteq M$ and $\Phi$ is $\mathcal{M}$-amenable since $\mathcal{N}$ is a faithful rank extension of $\mathcal{M}$. So intuitively speaking, $\Phi$ is the subset of the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) that consists of sentences whose constants are in $\{c\}\cup\{c_{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ need not be standard formulae. Also note that the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Now let:
$\Gamma:=\left\{t(c,c_{m})\in M:t(c,c_{m})\in\Phi\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,c_{m})>c_{\theta}\right)\in\Phi\right\},$
where $t$ is a definable function in the language $\mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{c_{m}:m\in M\}.$ So, officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,c_{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
(1) $\ \ \left[\exists!x\varphi(c,c_{m},x)\right]\in\Phi.$
(2) $\ \ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
(3) $\ \ \forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow c_{\theta}\in x\right)\right]\in\Phi.$
@@
$\Gamma\neq\varnothing\wedge\forall t\left(t\in\Gamma\longrightarrow\left[\exists t^{\prime}\in\Gamma\wedge\left((t^{\prime}\in t)\in\Phi\right)\right]\right).$
In contrast to the proof of Theorem 4.1 we cannot at this point conclude that there is a countable descending chain $\left\langle t_{n}:n\in\omega\right\rangle$ in $\mathcal{M}$ since $\mathrm{DC}$ need not hold in $\mathcal{M}$. Instead, we will use the following argument that takes advantage of the Reflection Theorem. Choose $k\in\omega$ such that $\psi$ is a $\Sigma_{k}(\Phi)$-statement, and use the Reflection Theorem in $\left(\mathcal{M},\Phi\right)$ to pick $\mu\in\mathrm{Ord}^{\mathcal{M}}$ such that:

### v6

@@ 3. $\mathbb{C}_{3}(n):$
This section contains the proof of Theorem B of the abstract (see Corollary 4.2). Recall from Definition 2.4 and Remark 2.5 that the notion of a faithful extension is a generalization of the notion of a conservative extension. The notion of a full satisfaction class was defined in Definition 2.7(b).
4.1. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathsf{ZFC}$ such that $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\mathrm{faith}}\mathcal{N}$, $\mathcal{N}$ fixes $\omega^{\mathcal{M}},$ and $\mathcal{N}$ is taller than $\mathcal{M}$. Then:
  1. (a)
    There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{M=N}_{\gamma}$, in which case $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$. [^26] As pointed out in Remark 4.5, both cases of the dichotomy can be realized.
  2. (b)
    In particular, there is a full satisfaction class $S$ for $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof [^27] The proof strategy of Theorem 4.1 is a variant of the proofs of the following results: [En-2, Theorem 1.5], [En-3, Theorem 3.3], [EH, Theorem 2.1], [En-6, Theorem 2.1.3], and [En-7, Lemma 2.19]. A variant of the same strategy is used in the proof of Theorem 4.5. The origins of the strategy can be traced to Kaufmann’s refinement [Kau, Lemma 1.4] of a Skolem hull argument due to Keisler and Silver [KS, Theorem 2.1].. Fix some $\mathcal{N}$-ordinal $\lambda\in\mathrm{Ord}^{\mathcal{N}}$ that dominates each $\mathcal{M}$-ordinal, and some $\mathcal{N}$-ordinal $\beta>\lambda$. Note that:
@@
$K_{m}:=\{a\in\mathrm{V}:\mathcal{N}\models a\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m)\},$
where $x\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in the structure $(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m).$
Note that the only allowable parameters used in a definition of $x$ are $\lambda$ and $m$. Thus, intuitively speaking, within $\mathcal{N}$ the set $K_{m}$ consists of elements $a$ of $\mathrm{V}_{\beta}$ such that $a$ is first order definable in $\left(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m\right)$. Also note that since we are not assuming that $\mathcal{M}$ is $\omega$-standard, an element $a\in K_{m}$ need not be definable in the structure $\left(\mathcal{N}_{\beta},\vartriangleleft,\lambda,m\right)$ in the real world. Next we move outside of $\mathcal{N}$ and define $K$ as follows:
@@
[…]
@@
[…]
@@
Clearly $O^{\ast}$ is nonempty since $\lambda\in O$ . We now consider the following two cases. As we shall see, Case I leads to the conclusion of the theorem, and Case II is impossible.
@@
[…]
@@
$\Phi_{m}:=\{x\in M:\mathcal{N}\models$ $x$ is (the code of) a formula $\varphi(c,\dot{m})$ such that $\varphi(\dot{\lambda},\dot{m})\in s_{\beta}\}\mathrm{.}$
So intuitively speaking, $\Phi$ is the result of replacing $\dot{\lambda}$ by $c$ (where $c$ is a fresh constant) in the sentences in the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) whose constants are in $\{\dot{\lambda}\}\cup\{\dot{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ might be nonstandard formulae.
@@
$\Gamma:=\left\{\begin{array}[]{c}t(c,\dot{m})\in M:t(c,\dot{m})\in\Phi,\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,\dot{m})>\dot{\theta}\right)\in\Phi,\ \mathrm{and}\\ t\ \mathrm{is\ a\ definable\ term\ in\ the\ language}\ \mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{\dot{m}:m\in M\}\end{array}\right\}.$
Officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,\dot{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
  1. $(i)$
    $\left[\exists!x\varphi(c,\dot{m},x)\right]\in\Phi.$
  2. $(ii)$
    $\left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
  3. $(iii)$
    $\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow\dot{\theta}\in x\right)\right]\in\Phi.$
@@ 3. $(iii)$
$\psi:=\forall t\left[t\in\Gamma\longrightarrow\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi\right].$
The veracity of the dependent choice scheme in $\mathcal{M}$ (see part (f) of Definition 2.1), together with the facts that $\Phi$ is $\mathcal{M}$-amenable and $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$ make it clear that there is a sequence $s=\left\langle t_{n}:n\in\omega^{\mathcal{M}}\right\rangle$ in $\mathcal{M}$ such that:
@@
$\mathcal{N}\models\left\langle t_{n}^{(\mathrm{V}_{\beta},\in)}(\dot{\lambda}):n\in\omega\right\rangle$ is an infinite decreasing sequence of ordinals.
In the above $t_{n}(\dot{\lambda})$ is the term obtained by replacing $c$ with $\dot{\lambda}$ in $t_{n},$ and $t_{n}^{(\mathrm{V}_{\beta},\in)}(\dot{\lambda})$ is the interpretation of $t_{n}(\dot{\lambda})$ in $(\mathrm{V},\in)$ using the Tarskian satisfaction class for $(\mathrm{V}_{\beta},\in)$. $\square$
4.2. Corollary. If $\mathcal{M}\models\mathsf{ZFC}$, $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ fixes $\omega^{\mathcal{M}},$ then $\mathcal{M}$ is cofinal in $\mathcal{N}$. Thus a conservative elementary extension of $\mathcal{M}$ that fixes $\omega^{\mathcal{M}}$ is a cofinal extension.
Proof. If $\mathcal{M}\models\mathsf{ZFC}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ is not cofinal in $\mathcal{N}$, then $\mathcal{N}$ is taller than $\mathcal{M}$. But then by part (b) of Theorem 4.1, there is an $\mathcal{M}$-definable full satisfaction class over $\mathcal{M}$, which contradicts Tarski’s Undefinability of Truth Theorem. $\square$
In Theorem 4.4 below we show that the conclusion of Theorem 4.1 holds if $\mathsf{ZFC}$ is weakened to $\mathsf{ZF}\mathrm{,}$ but the assumption that $\mathcal{N}$ fixes $\omega^{\mathcal{M}}$ is strengthened to $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}$. We first establish a lemma that will come handy in the proof of Theorem 4.4.
4.3. Lemma Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then $\mathcal{M}\subsetneq_{\mathrm{rank}}\mathcal{N}$.
@@
[…]
@@
(3) $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$ and $\mathrm{Ext}_{\mathcal{M}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b^{\prime}).$
So by (1) and (2),
(4) $\mathrm{Ext}_{\mathcal{N}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{N}}(a).$
@@
which by Extensionality makes it clear that $b^{\prime}=b.$ $\square$
  • The proof of Theorem 4.4 below is a variant of the proof of Theorem 4.1, and will strike the reader who has worked through the proof of Theorem 4.1 as repetitious, but the proof is a bit more involved here since we do not have access to the axiom of choice in $\mathcal{M}$ or in $\mathcal{N}$. Note that in Theorem 4.4 $\mathcal{N}$ is assumed to end extend $\mathcal{M}$, whereas in Theorem 4.1 $\mathcal{N}$ is only assumed to fix $\omega^{\mathcal{M}}.$ Of course Theorem 4.4 follows from Theorem 4.1 if $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZFC}$.
4.4. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then:
  1. (a)
@@
  2. (b)
    In particular, there is a full satisfaction class $S$ for $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
[…]
@@
where $x\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in $(\mathrm{V}_{\beta},\in)$ using at most the parameters $\lambda$ and $m.$
@@
[…]
  1. (I)
    $O^{\ast}$ has a least element (under $\in^{\mathcal{N}}).$
  2. (II)
    $O^{\ast}$ has no least element.
Case I. Let $\gamma_{0}$ be the least element of $O^{\ast}.$ It is easy to see that $\gamma_{0}$ is a limit ordinal in $\mathcal{N}$. We claim that:
@@
To verify the claim, by the Tarski criterion of elementarity, it suffices to show that if $\mathcal{N}_{\gamma_{0}}\models\exists y\varphi(y,m)$ for some $m\in M$, then there is some $m_{0}\in M$ such that $\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$ Let $\gamma_{m}$ be the first ordinal below $\gamma_{0}$ such that:
@@
So $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameters $\gamma_{0}$ and $m$, and since $\gamma_{0}\in O$ by assumption, this shows that $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameter $m.$ So $\gamma_{m}\in O.$ By the fact that $\gamma_{0}$ was chosen to be the least element of $O^{\ast}$, this shows that $\gamma_{m}\in\mathrm{Ord}^{\mathcal{M}},$ which makes it clear that there is some $m_{0}\in M$ such that:
@@
This concludes the verification of $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. To finish the proof of Theorem 4.4 it suffices to show that Case II is impossible. Suppose to the contrary that $O^{\ast}$ has no least element. Within $\mathcal{N}$ let $s_{\beta}$ be the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$. For each $m\in M$ we define:
$\Phi_{m}:=\{\varphi(c,\dot{m})\in M:\mathcal{N}\models$ $\varphi(\dot{\lambda},\dot{m})\in s_{\beta}\}.$
Note the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Next we define:
@@
Observe that $\Phi\subseteq M$ and $\Phi$ is $\mathcal{M}$-amenable since $\mathcal{N}$ is a faithful rank extension of $\mathcal{M}$. So intuitively speaking, $\Phi$ is the subset of the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) that consists of sentences whose constants are in $\{c\}\cup\{\dot{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ need not be standard formulae. Also note that the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Now let:
$\Gamma:=\left\{\begin{array}[]{c}t(c,\dot{m})\in M:t(c,\dot{m})\in\Phi,\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,\dot{m})>\dot{\theta}\right)\in\Phi,\ \mathrm{and}\\ t\ \mathrm{is\ a\ definable\ term\ in\ the\ language}\ \mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{\dot{m}:m\in M\}\end{array}\right\},$
So, officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,\dot{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
  1. $(i)$
    $\left[\exists!x\varphi(c,\dot{m},x)\right]\in\Phi.$
  2. $(ii)$
    $\left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
  3. $(iii)$
    $\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow\dot{\theta}\in x\right)\right]\in\Phi.$
@@
$\Gamma\neq\varnothing\wedge\forall t\left[t\in\Gamma\longrightarrow\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi\right].$
In contrast to the proof of Theorem 4.1 we cannot at this point conclude that there is a countable descending chain $\left\langle t_{n}:n\in\omega\right\rangle$ in $\mathcal{M}$ since $\mathsf{DC}$ need not hold in $\mathcal{M}$. Instead, we will use the following argument that takes advantage of the Reflection Theorem. Choose $k\in\omega$ such that $\psi$ is a $\Sigma_{k}(\Phi)$-statement, and use the Reflection Theorem in $\left(\mathcal{M},\Phi\right)$ to pick $\mu\in\mathrm{Ord}^{\mathcal{M}}$ such that: